# FlowCast: Advancing Precipitation Nowcasting with Conditional Flow Matching

### v1

@@ 5 Conclusion
In this paper, we introduced FlowCast, the first model to apply Conditional Flow Matching (CFM) to precipitation nowcasting. Our experiments on the SEVIR and ARSO datasets show that FlowCast achieves state-of-the-art performance. Through direct ablation studies, we showed that the CFM objective is not only more accurate than a traditional diffusion objective on the same architecture but also vastly more efficient. FlowCast maintains high forecast quality with as few as a single sampling step, a regime where diffusion models fail. Our results firmly establish CFM as a powerful, efficient, and practical alternative for high-dimensional spatiotemporal forecasting.
@@ A.1.1 VAE
The VAE was configured with the following hyperparameters:
@@ A.1.2 FlowCast
We conducted an ablation study to compare various ODE solvers, including adaptive methods (Adaptive Heun, Dormand-Prince 5) and fixed-step methods (Euler, Midpoint, Runge-Kutta 4). Since no significant performance differences were observed, we selected the Euler method with 10 steps for its computational efficiency and simplicity.

### v2

@@ 5 Conclusion
In this paper, we introduced FlowCast, the first fully probabilistic model applying Conditional Flow Matching (CFM) as a direct noise-to-data generative framework for precipitation nowcasting. Our experiments on the SEVIR and ARSO datasets show that FlowCast achieves state-of-the-art performance. Through direct ablation studies, we showed that the CFM objective is not only more accurate than a traditional diffusion objective on the same architecture but also vastly more efficient. FlowCast maintains high forecast quality with as few as a single sampling step, a regime where diffusion models fail. Our results firmly establish CFM as a powerful, efficient, and practical alternative for high-dimensional spatiotemporal forecasting.
@@ A.1.1 VAE
The VAE was configured with the hyperparameters shown in Table 7. Different warmup periods were empirically selected based on the convergence speed of the VAE on the respective datasets; ARSO converged faster due to less diversity in the data, attributed to factors such as its fixed geographical coverage and the significantly smaller stride used during sequence extraction.
@@
Data Preprocessing and Padding.
To accommodate the VAE’s downsampling factor of $f=8$, inputs must be spatially divisible by the downsampling rate. For the ARSO dataset, the native resolution of $301\times 401$ is not divisible by 8. We handle this by applying replication padding to the input frames to reach the nearest multiple of 16 prior to encoding. Specifically, the height is padded from 301 to 304, and the width from 401 to 416. This results in the latent dimensions of $38\times 52$ reported in Table 2 ($304/8$ and $416/8$, respectively). During inference, the generated fields are cropped back to the original $301\times 401$ dimensions before evaluation metrics are computed. SEVIR dimensions ($384\times 384$) are naturally divisible by 8, requiring no padding.
@@ A.1.2 FlowCast
We conducted an ablation study to compare various ODE solvers, including adaptive methods (Adaptive Heun, Dormand-Prince 5) and fixed-step methods (Euler, Midpoint, Runge-Kutta 4) on the first 10% of the SEVIR test set using a single NVIDIA A100 GPU. For adaptive solvers, a relative and absolute tolerance of $10^{-2}$ and $10^{-3}$ were used, respectively. Since no significant performance differences were observed, as shown in Table 9, we selected the Euler method with 10 steps for its computational efficiency and simplicity.
Table 9: Ablation study: ODE solvers. Results highlight minor differences in performance between the different solvers.
Solver CRPS $\downarrow$ CSI-M $\uparrow$ CSI-P16-M $\uparrow$ FSS-M-P16 $\uparrow$ HSS-M $\uparrow$ FAR-M $\downarrow$ Time/Seq. (s)
Euler (1 steps) 0.0207 0.454 0.504 0.763 0.571 0.337 2.6
Euler (10 steps) 0.0168 0.455 0.514 0.764 0.572 0.338 24
Dormand-Prince 5 0.0168 0.450 0.516 0.762 0.567 0.341 46
Midpoint (10 steps) 0.0167 0.451 0.516 0.762 0.568 0.341 44
Runge-Kutta 4 (10 steps) 0.0167 0.451 0.516 0.762 0.567 0.341 83
Adaptive Heun 0.0167 0.451 0.516 0.762 0.568 0.341 50
@@ A.1.3 Evaluation Metrics
Fractions Skill Score (FSS):
We compute FSS to evaluate spatial alignment. Let $S_{f}(i,j)$ and $S_{o}(i,j)$ denote the fraction of pixels exceeding the threshold within a neighborhood of size $n\times n$ centered at $(i,j)$ for the forecast and observation, respectively. The FSS is calculated as:
$\mathrm{FSS}=1-\frac{\sum_{i,j}[S_{f}(i,j)-S_{o}(i,j)]^{2}}{\sum_{i,j}S_{f}(i,j)^{2}+\sum_{i,j}S_{o}(i,j)^{2}}$ (5)